# Can RLHF be More Efficient with Imperfect Reward Models? A Policy Coverage Perspective

### v1

@@ 4.2 Main Results and Interpretation
We decompose the total regret into two parts depending on their origins. $\text{Reg}_{\texttt{OL}}^{(t)}$ comes from the regret by running the online algorithm $\texttt{Alg}_{\texttt{OL}}$. It is weighted by $\alpha$ since we only allocate $\alpha$-proportion of the samples for $\texttt{Alg}_{\texttt{OL}}$. $\text{Reg}_{\texttt{Trf}}^{(t)}$ represents the regret from transfer policies. The first term in Eq. (8) reflects the benefits of utilizing transfer policy over online learning. Here $\Delta_{\min}$ is contributed by source reward models $\{r^{w}\}_{w\in[W]}$, and the term $\iota^{k(\tau),n(\tau)}$ is due to the “self-transfer policy” $\pi_{\texttt{OFF}}$, as we derived in Lem. 4.2. The second term in Eq. (8) results from the imperfection of source reward models: without prior knowledge on their quality, additional cost has to be paid during exploration.
@@
Remarks on the choice of $\alpha$  Our algorithm treats $\alpha$ as a hyperparameter. Depending on the concrete scenarios, if we have stronger prior beliefs that at least one of $\{r^{w}\}_{w\in[W]}$ is similar to $r^{*}$, we should prefer larger $\alpha$. Besides, one may gradually decay $\alpha$ to 0 when the iteration number is large enough. As a result, the transfer regret in Eq. (8) dominates, leading to the total regret growing with $\widetilde{O}(\sqrt{T})$ over time.
Improved regret bound in online RLHF Although our focus is transfer learning, our results can be extended to the standard online RLHF setting when no source tasks are present (i.e., $W=0$). The corollary below shows that a variant of TPO only utilizing self-transfer learning can achieve an $\widetilde{O}(\sqrt{T})$ online regret, thereby strictly improving existing results (Xiong et al., 2024; Xie et al., 2024; Cen et al., 2024; Zhang et al., 2024).
Corollary 4.4.
Suppose $\texttt{Alg}_{\texttt{OL}}$ is a no-regret instance satisfying Def. D.2, then w.p. $1-\delta$, running Alg. 1 with $W=0$ and $T_{0}=\widetilde{\Theta}(\kappa^{2}(e^{\frac{2R}{\beta}})\mathcal{C}(\Pi)\beta^{-2})$, by choosing $\alpha=1$ for $t\leq T_{0}$ and $\alpha=0$ for $T_{0}<t\leq T$, it achieves an online regret $\widetilde{O}(e^{2R}\sqrt{T})$.
5 From Theory to an Empirical Algorithm
@@
$\widehat{\text{WR}}^{\pi^{k}_{\texttt{OL}}}:={\mathbb{P}}_{r^{*}}(\pi^{k}_{\texttt{OL}}\succ\pi^{k}_{\texttt{OL}})=0.5.$
@@
Collect data $(s^{k,n},a^{k,n},\widetilde{a}^{k,n},y^{k,n})\sim\rho\times\pi^{k,n}\times\pi^{k}_{\texttt{OL}}\times{\mathbb{P}}_{r^{*}}(\cdot|\cdot,\cdot,\cdot)$.
@@
      $\pi^{k+1}_{\texttt{OL}}\leftarrow{\texttt{DPO}}(\pi^{k}_{\texttt{OL}},\mathcal{D}^{k,N+1})$;
@@
Note that we may not identify the policy with the best coverage for $\pi^{*}_{r^{*}}$ through the lower bound above. However, it still provides useful guidance for practice: we can filter out policies yielding high lower bound. In Lem. 5.1, for any fixed $\gamma$ and comparator $\bar{\pi}$, the lower bound for ${\texttt{Cov}}^{\pi^{*}_{r^{*}}|\pi}$ increases as ${\mathbb{P}}_{r^{*}}(\pi\succ\bar{\pi})$ decay to 0, which suggests prioritizing transferring from policies with high win rates.
@@
From Insights to Practice  Next, we walk through empirical TPO in Alg. 3 and explain we integrate these insights into the algorithm design. Alg. 3 builds upon the Iterative-DPO framework (Xiong et al., 2024; Yuan et al., 2024), which iteratively collects data and optimize the policy using the DPO (Rafailov et al., 2024) loss. We start by initializing the online learning policy $\pi^{1}_{\texttt{OL}}$ with reference policy $\pi_{\texttt{ref}}$. For computational efficiency, in each iteration $k$, we avoid separately computing online and offline policies as done in TPO. Instead, we optimize the DPO loss once to update the online policy $\pi^{k}_{\texttt{OL}}$. During data collection, the algorithm selects the policy $\pi^{k,n}\in\{\pi^{*}_{r^{*}}\}_{w=1}^{W}\cup\{\pi^{k}_{\texttt{OL}}\}$ with the highest win rate when competing against $\pi^{k}_{\texttt{OL}}$. Intuitively, we encourage transfer learning if $\{\pi^{*}_{r^{*}}\}_{w=1}^{W}$ includes high-quality candidates; otherwise, the algorithm conducts standard online DPO by default. This strategy also aligns with the heuristic principle: learn from an expert until surpassing it. Lastly, since the win rates are unknown in advance, the selection process is formulated as a multi-armed bandit problem. We employ a UCB subroutine (line 3) to balance the exploration and exploitation during the win rates estimation.
@@ 6 Experiments
In this section, we evaluate Alg. 3 on the summarization task using the XSum dataset (Narayan et al., 2018). We consider T5 series models (Raffel et al., 2020) and choose T5-small (80M) as the base model for fine-tuning. Human reward $r^{*}$ is simulated by the reward model (Dong et al., 2024) distilled from Llama3-8B (Dubey et al., 2024). For the source rewards in transfer learning, we consider a collection of imperfect reward models and LLM policies, including, (a) ROUGE-Lsum score (Lin, 2004), (b) BERTScore (Zhang et al., 2019), (c) T5-base (250M), (d) T5-large (770M). As illustrated in Sec. 2.2, for LLM policies (c) and (d), we treat their log-probability predictions on the given prompt $s$ and response $a$ as the reward scores. To save space, we present main results below and defer the concrete setups and a few additional experiment results to Appx. I.
@@
Experiment Results and Interpretation  We run Alg. 3 for $K=3$ iterations and compare its performance with three baselines: (I) vanilla iterative-DPO without transfer learning; (II) purely exploiting the worst source reward; (III) purely exploiting the best source reward. Concretely, baseline (I) removes the transfer learning component in Alg. 3 by assigning $\pi^{k,n}=\pi_{\texttt{OL}}^{k}$ for all $n\in[N]$. For baselines (II) and (III), the worst (ROUGE-LSum) and best (T5-Large) reward models from the candidates (a)-(d) are selected, and pure transfer learning is then performed using the responses recommended by the chosen reward models, i.e., $\pi^{k,n}=\pi^{*}_{r^{w}}$ in Alg. 3 for the selected $r^{w}$. Here the worst and best reward models are selected based on the final policy value when aligning with the given reward model.
@@ Appendix C Offline Learning Results in Previous Literature
$\forall\pi\in\text{conv}(\Pi),\quad J_{\beta}(\pi)-J_{\beta}(\pi_{\texttt{OFF}})\leq C_{\texttt{OFF}}e^{2{R}}\cdot{\texttt{Cov}}^{\pi|\pi_{\text{mix}}^{\mathcal{D}}}\sqrt{\frac{1}{|\mathcal{D}|}\log\frac{|\Pi|}{\delta}},$
@@
Next, we adapt the proofs in (Liu et al., 2024). Note that we can directly start with their Eq. (D.4), because their bounds in Eq. (D.2) and Eq. (D.3) only involve optimality of the choice of $\pi_{\texttt{OFF}}$ and realizability. We move the KL-regularization terms to the LHS and merge to $J_{\beta}(\pi)$ and $J_{\beta}(\pi_{\texttt{OFF}})$, and we choose $\pi_{\texttt{ref}}$ as the base policy in RPO. The adapted results to our notations would be:
$\forall\pi\in\text{conv}(\Pi),~{}J_{\beta}(\pi)$ $-J_{\beta}(\pi_{\texttt{OFF}})$
@@ Appendix E Proofs for Results in Section 3
Lem. 3.1 then holds directly as a corollary by combining with Lem. H.1.
@@ E.2 Another Bound for Policy Coverage Coefficient
$\quad J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{\texttt{OFF}})\leq C_{\texttt{OFF}}e^{2{R}}\cdot{\texttt{Cov}}^{\pi^{*}_{r^{*}}|\pi_{\text{mix}}^{T}}\sqrt{\frac{1}{T}\log\frac{|\Pi|}{\delta}},$
@@
Under Assump. 2.1, suppose $\pi^{1},...,\pi^{T}$ is generated by a no-regret online learning algorithm with $\sum_{t=1}^{T}J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi^{t})=\widetilde{O}(\mathcal{C}(\Pi)\sqrt{T})$ for some structural complexity measure $\mathcal{C}(\Pi)$, as long as $T=\widetilde{\Omega}(\beta^{-2}\mathcal{C}(\Pi)^{2}\kappa^{2}(e^{\frac{2{R}}{\beta}}))$, running RPO yields an offline policy s.t. $J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{\texttt{OFF}})=\widetilde{O}(e^{2{R}}T^{-\frac{1}{2}})$.
@@ F.2 Some Useful Lemmas
$\widehat{V}(\pi_{\texttt{OFF}};\mathcal{D})\leq$ $J_{\beta}(\pi_{\texttt{OFF}})-J_{\beta}(\pi_{\texttt{ref}})$
$\widehat{V}(\pi_{\texttt{OFF}};\mathcal{D})\geq$ $J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{\texttt{ref}})-c^{\prime}\cdot{R}e^{2{R}}\cdot\Big{(}{\texttt{Cov}}^{\pi^{*}_{r^{*}}|\pi_{\text{mix}}^{\mathcal{D}}}\wedge\frac{\sqrt{\mathcal{C}(\Pi)}}{\alpha}\Big{)}\cdot\sqrt{\frac{1}{|\mathcal{D}|}\log^{c_{0}}\frac{|\Pi|T}{\delta}},$
@@
$\widehat{V}(\pi_{\texttt{OFF}};\mathcal{D}):=$ $\mathbb{E}_{\rho,\pi_{\texttt{OFF}}}[{\widehat{r}}_{\texttt{OFF}}]-\mathbb{E}_{\rho,\pi_{\texttt{ref}}}[{\widehat{r}}_{\texttt{OFF}}]-\beta\text{KL}(\pi_{\texttt{OFF}}\|\pi_{\texttt{ref}})$
$+\frac{1}{\eta}L_{\mathcal{D}}({\widehat{r}}_{\texttt{OFF}})-\frac{1}{\eta}L_{\mathcal{D}}({\widehat{r}}_{\texttt{MLE}})-\texttt{bonus},$
@@
$\widehat{V}(\pi_{\texttt{OFF}};\mathcal{D})$
$\leq$ $\mathbb{E}_{\rho,\pi_{\texttt{OFF}}}[r^{*}]-\mathbb{E}_{\rho,\pi_{\texttt{ref}}}[r^{*}]-\beta\text{KL}(\pi_{\texttt{OFF}}\|\pi_{\texttt{ref}})+\frac{1}{\eta}L_{\mathcal{D}}(r^{*})-\frac{1}{\eta}L_{\mathcal{D}}({\widehat{r}}_{\texttt{MLE}})-\texttt{bonus}$ (Pessimistic estimation of ${\widehat{r}}_{\texttt{OFF}}$ in Eq. (12))
$\leq$ $J_{\beta}(\pi_{\texttt{OFF}})-J_{\beta}(\pi_{\texttt{ref}})+\frac{2}{\eta|\mathcal{D}|}\log\frac{|\Pi|}{\delta}-\texttt{bonus}$ (Lem. H.2)
$\leq$ $J_{\beta}(\pi_{\texttt{OFF}})-J_{\beta}(\pi_{\texttt{ref}})+2c\cdot e^{2{R}}\sqrt{\frac{1}{|\mathcal{D}|}\log\frac{|\Pi|T}{\delta}}-\texttt{bonus}.$ (14)
@@
$J_{\beta}(\pi)-J_{\beta}(\pi_{\texttt{ref}})-\widehat{V}(\pi_{\texttt{OFF}};\mathcal{D})$
@@
$-\Big{(}\mathbb{E}_{\rho,\pi_{\texttt{OFF}}}[{\widehat{r}}_{\texttt{OFF}}]-\mathbb{E}_{\rho,\pi_{\texttt{ref}}}[{\widehat{r}}_{\texttt{OFF}}]-\beta\text{KL}(\pi_{\texttt{OFF}}\|\pi_{\texttt{ref}})+\frac{1}{\eta}L_{\mathcal{D}}({\widehat{r}}_{\texttt{OFF}})\Big{)}+\frac{1}{\eta}L_{\mathcal{D}}({\widehat{r}}_{\texttt{MLE}})+\texttt{bonus}$
$\leq$ $\Big{(}\mathbb{E}_{\rho,\pi}[r^{*}]-\mathbb{E}_{\rho,\pi_{\texttt{ref}}}[r^{*}]-\beta\text{KL}(\pi\|\pi_{\texttt{ref}})\Big{)}-\min_{r\in\mathcal{R}^{\Pi}}\Big{(}\mathbb{E}_{\rho,\pi}[r]-\mathbb{E}_{\rho,\pi_{\texttt{ref}}}[r]-\beta\text{KL}(\pi\|\pi_{\texttt{ref}})+\frac{1}{\eta}L_{\mathcal{D}}(r)\Big{)}$ (Optimality of $\pi_{\texttt{OFF}}$ in RPO;)
@@
$J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{\texttt{ref}})-\widehat{V}(\pi_{\texttt{OFF}};\mathcal{D})\leq 4c_{2}\cdot e^{2{R}}\cdot{\texttt{Cov}}^{\pi^{*}_{r^{*}}|\pi_{\text{mix}}^{\mathcal{D}}}\cdot\sqrt{\frac{1}{|\mathcal{D}|}\log\frac{|\Pi|T}{\delta}}+\texttt{bonus}.$ (16)
@@
$J_{\beta}(\pi^{k-1}_{\text{mix}})-J_{\beta}(\pi_{\texttt{ref}})-\widehat{V}(\pi_{\texttt{OFF}};\mathcal{D})\leq 4c_{2}\cdot e^{2{R}}\cdot\frac{2}{\alpha}\cdot\sqrt{\frac{1}{|\mathcal{D}|}\log\frac{|\Pi|T}{\delta}}.$
@@
$J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{\texttt{ref}})-\widehat{V}(\pi_{\texttt{OFF}};\mathcal{D})$
$\leq$ $J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{{\text{mix}}}^{k-1})+J_{\beta}(\pi_{{\text{mix}}}^{k-1})-J_{\beta}(\pi_{\texttt{ref}})-\widehat{V}(\pi_{\texttt{OFF}};\mathcal{D})$
@@
$\widehat{V}(\pi_{\texttt{OFF}};\mathcal{D})\leq$ $J_{\beta}(\pi_{\texttt{OFF}})-J_{\beta}(\pi_{\texttt{ref}})$
$\widehat{V}(\pi_{\texttt{OFF}};\mathcal{D})\geq$ $J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{\texttt{ref}})-c^{\prime}{R}e^{2{R}}\cdot\Big{(}{\texttt{Cov}}^{\pi^{*}_{r^{*}}|\pi_{\text{mix}}^{\mathcal{D}}}\wedge\frac{\sqrt{\mathcal{C}(\Pi)}}{\alpha}\Big{)}\cdot\sqrt{\frac{1}{|\mathcal{D}|}\log^{c_{0}}\frac{|\Pi|T}{\delta}}.$

### v2

@@ 4.2 Main Results and Interpretation
We decompose the total regret into two parts depending on their origins. $\text{Reg}_{\texttt{OL}}^{(t)}$ comes from the regret by running the online algorithm $\texttt{Alg}_{\texttt{OL}}$. It is weighted by $\alpha$ since we only allocate $\alpha$-proportion of the samples for $\texttt{Alg}_{\texttt{OL}}$. $\text{Reg}_{\texttt{Trf}}^{(t)}$ represents the regret from transfer policies. The first term in Eq. (8) reflects the benefits of utilizing transfer policy over online learning. Here $\Delta_{\min}$ is contributed by source reward models $\{r^{w}\}_{w\in[W]}$, and the term $\iota^{k(\tau),n(\tau)}$ is due to the “self-transfer policy” $\pi_{\text{Dstl}}$, as we derived in Lem. 4.2. The second term in Eq. (8) results from the imperfection of source reward models: without prior knowledge on their quality, additional cost has to be paid during exploration.
@@
$\widehat{\text{WR}}^{\pi^{k}_{\texttt{OL}}}\leftarrow{\mathbb{P}}_{r^{*}}(\pi^{k}_{\texttt{OL}}\succ\pi^{k}_{\texttt{OL}})=0.5.$ // $\widehat{\text{WR}}^{\pi^{k}_{\texttt{OL}}}$ can be treated as a hyperparameter taking value other than 0.5.
@@
Collect online data $(s^{k,n},a^{k,n},\widetilde{a}^{k,n},y^{k,n})\sim\rho\times\pi^{k,n}\times\pi^{k}_{\texttt{OL}}\times{\mathbb{P}}_{r^{*}}(\cdot|\cdot,\cdot,\cdot)$.
@@
      $\pi^{k+1}_{\texttt{OL}}\leftarrow\text{Alg}_{\text{PO}}(\pi^{k}_{\texttt{OL}},\mathcal{D}^{k,N+1})$;
@@
Choice of $\alpha$ and the total regret of TPO  Our algorithm treats $\alpha$ as a hyperparameter. Based on the discussion above, with proper choice of $\alpha$ we have the following corollary.
Corollary 4.4 (Total Regret of TPO).
By choosing XPO (Xie et al., 2024) as $\texttt{Alg}_{\texttt{OL}}$ and setting $\alpha=e^{-\frac{R}{\beta}}\leq{\texttt{Cov}}_{\infty}^{-1}(\Pi)$, Thm. 4.3 implies that TPO achieves $\widetilde{O}(W\sqrt{T})$ regret if $T<\frac{W^{2}}{\Delta^{2}_{\min}}$, or $\widetilde{O}(\sqrt{T})$ regret if $T>\frac{W^{2}}{\Delta^{2}_{\min}}$ and large enough.
Depending on the concrete scenarios, if we have stronger prior beliefs that one of $\{r^{w}\}_{w\in[W]}$ is similar to $r^{*}$, we should prefer larger $\alpha$. Besides, one may gradually decay $\alpha$ to 0 when the iteration number is large enough, which can also result in a total regret growing with $\widetilde{O}(\sqrt{T})$ over time.
Improved regret bound in standard online RLHF without source rewards Although our focus is transfer learning, our results can be extended to the standard online RLHF setting where no source tasks are present (i.e., $W=0$). As implied by Corollary 4.4, only utilizing self-transfer learning can result in an $\widetilde{O}(\sqrt{T})$ online regret, thereby strictly improving existing results (Xiong et al., 2024; Xie et al., 2024; Cen et al., 2024; Zhang et al., 2024).
5 From Theory to an Empirical Algorithm
@@
Note that we may not identify the policy with the best coverage for $\pi^{*}_{r^{*}}$ through the lower bound above. However, it still provides useful guidance for practice: we can filter out policies yielding high lower bound. In Lem. 5.1, for any fixed $\gamma$ and comparator $\bar{\pi}$, the lower bound for ${\texttt{Cov}}^{\pi^{*}_{r^{*}}|\pi}$ increases as ${\mathbb{P}}_{r^{*}}(\pi\succ\bar{\pi})$ decay to 0, suggesting prioritizing transferring from policies with high win rates.
@@
From Insights to Practice  Next, we walk through empirical TPO in Alg. 3 and explain how we integrate these insights into the algorithm design. Alg. 3 utilizes an iterative online learning framework, which repeatedly collects online data and optimizes the policy. We start by initializing the online learning policy $\pi^{1}_{\texttt{OL}}$ with the reference policy $\pi_{\texttt{ref}}$. For computational efficiency, in each iteration $k$, we avoid separately computing online exploration policies and self-transfer learning policies as done in TPO. Instead, we only compute one policy $\pi^{k}_{\texttt{OL}}$ (updated from $\pi^{k-1}_{\texttt{OL}}$) by $\text{Alg}_{\text{PO}}$. Here $\text{Alg}_{\text{PO}}$ is a placeholder for an arbitrary Policy Optimization algorithm, and we do not restrict the concrete choice. Such a design increases the modularity of our empirical TPO, making it possible to combine with various policy optimization methods and enhance their performance. For example, $\text{Alg}_{\text{PO}}$ may be instantiated by DPO (Rafailov et al., 2024), resulting in a transfer learning framework built on iterative-DPO (Xiong et al., 2024; Yuan et al., 2024). Besides, one may consider other advanced (online) methods, such as XPO (Xie et al., 2024), IPO (Azar et al., 2024), etc.
As the core ingredients of our empirical TPO, during data collection, the algorithm selects the policy $\pi^{k,n}\in\{\pi^{*}_{r^{*}}\}_{w=1}^{W}\cup\{\pi^{k}_{\texttt{OL}}\}$ with the highest win rate when competing against $\pi^{k}_{\texttt{OL}}$. Intuitively, we encourage transfer learning if $\{\pi^{*}_{r^{*}}\}_{w=1}^{W}$ includes high-quality candidates; otherwise, the algorithm conducts standard iterative policy optimization with $\text{Alg}_{\text{PO}}$ by default. This strategy also aligns with the heuristic principle: learn from an expert until surpassing it. Lastly, since the win rates are unknown in advance, the selection process is formulated as a multi-armed bandit problem. We employ a UCB subroutine (line 3) to balance the exploration and exploitation during the win rates estimation.
@@ 6 Experiments
In this section, we evaluate Alg. 3 on the summarization task using the XSum dataset (Narayan et al., 2018). We consider T5 series models (Raffel et al., 2020) and choose T5-small (80M) as the base model for fine-tuning. Human reward $r^{*}$ is simulated by the reward model (Dong et al., 2024) distilled from Llama3-8B (Dubey et al., 2024). For the source rewards in transfer learning, we consider a collection of imperfect reward models and LLM policies, including, (a) ROUGE-Lsum score (Lin, 2004), (b) BERTScore (Zhang et al., 2019), (c) T5-base (250M), (d) T5-large (770M). As illustrated in Sec. 2.2, for LLM policies (c) and (d), we treat their log-probability predictions on the given prompt $s$ and response $a$ as the reward scores.
For $\text{Alg}_{\text{PO}}$, we consider three instantiations: DPO (Rafailov et al., 2024), IPO (Azar et al., 2024) and XPO (Xie et al., 2024). To save space, we present and interpret the results with DPO as the choice below and defer other experiment results and also the concrete setups to Appx. I.
Experiment Results and Interpretation  We run Alg. 3 for $K=3$ iterations and compare its performance with three baselines: (I) vanilla iterative-DPO without transfer learning (i.e., setting $\text{Alg}_{\text{PO}}$ = DPO and $W=0$); (II) purely exploiting the worst source reward—ROUGE score; (III) purely exploiting the best source reward—T5-large. Concretely, baseline (I) removes the transfer learning component in Alg. 3 by assigning $\pi^{k,n}=\pi_{\texttt{OL}}^{k}$ for all $n\in[N]$. For baselines (II) and (III), the worst (ROUGE-LSum) and best (T5-Large) reward models from the candidates (a)-(d) are selected, and pure transfer learning is then performed using the responses recommended by the chosen reward models, i.e., $\pi^{k,n}=\pi^{*}_{r^{w}}$ in Alg. 3 for the selected $r^{w}$. Here the worst and best reward models are selected based on the final policy value when aligning with the given reward model.
Notably, as suggested by the additional results in Appx. I, $\pi^{k}_{\texttt{OL}}$ improves over time and ${\mathbb{P}}_{r^{*}}(\pi^{*}_{r^{w}}\succ\pi^{k}_{\texttt{OL}})$ for any $w\in[W]$ continuously decreases. In iteration 3, our empirical TPO automatically switches back to online learning and avoids being restricted by the sub-optimality of source reward models. In the end, it results in higher win rates than purely exploiting the best source reward model T5-Large over 3 iterations.
@@ Appendix C Offline Learning Results in Previous Literature
$\forall\pi\in\text{conv}(\Pi),\quad J_{\beta}(\pi)-J_{\beta}(\pi_{\text{Dstl}})\leq C_{\texttt{OFF}}e^{2{R}}\cdot{\texttt{Cov}}^{\pi|\pi_{\text{mix}}^{\mathcal{D}}}\sqrt{\frac{1}{|\mathcal{D}|}\log\frac{|\Pi|}{\delta}},$
@@
Next, we adapt the proofs in (Liu et al., 2024). Note that we can directly start with their Eq. (D.4), because their bounds in Eq. (D.2) and Eq. (D.3) only involve optimality of the choice of $\pi_{\text{Dstl}}$ and realizability. We move the KL-regularization terms to the LHS and merge to $J_{\beta}(\pi)$ and $J_{\beta}(\pi_{\text{Dstl}})$, and we choose $\pi_{\texttt{ref}}$ as the base policy in RPO. The adapted results to our notations would be:
$\forall\pi\in\text{conv}(\Pi),~{}J_{\beta}(\pi)$ $-J_{\beta}(\pi_{\text{Dstl}})$
@@ Appendix E Proofs for Results in Section 3
Lem. 3.1 then holds directly as a corollary by combining with Lem. H.1, Lem. B.1 and the fact that $r^{w}\in[0,R]$ for all $w\in[W]$.
@@ E.2 Another Bound for Policy Coverage Coefficient
$\quad J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{\text{Dstl}})\leq C_{\texttt{OFF}}e^{2{R}}\cdot{\texttt{Cov}}^{\pi^{*}_{r^{*}}|\pi_{\text{mix}}^{T}}\sqrt{\frac{1}{T}\log\frac{|\Pi|}{\delta}},$
@@
Under Assump. 2.1, suppose $\pi^{1},...,\pi^{T}$ is generated by a no-regret online learning algorithm with $\sum_{t=1}^{T}J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi^{t})=\widetilde{O}(\mathcal{C}(\Pi)\sqrt{T})$ for some structural complexity measure $\mathcal{C}(\Pi)$, as long as $T=\widetilde{\Omega}(\beta^{-2}\mathcal{C}(\Pi)^{2}\kappa^{2}(e^{\frac{2{R}}{\beta}}))$, running RPO yields an offline policy s.t. $J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{\text{Dstl}})=\widetilde{O}(e^{2{R}}T^{-\frac{1}{2}})$.
@@ F.2 Some Useful Lemmas
$\widehat{V}(\pi_{\text{Dstl}};\mathcal{D})\leq$ $J_{\beta}(\pi_{\text{Dstl}})-J_{\beta}(\pi_{\texttt{ref}})$
$\widehat{V}(\pi_{\text{Dstl}};\mathcal{D})\geq$ $J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{\texttt{ref}})-c^{\prime}\cdot{R}e^{2{R}}\cdot\Big{(}{\texttt{Cov}}^{\pi^{*}_{r^{*}}|\pi_{\text{mix}}^{\mathcal{D}}}\wedge\frac{\sqrt{\mathcal{C}(\Pi)}}{\alpha}\Big{)}\cdot\sqrt{\frac{1}{|\mathcal{D}|}\log^{c_{0}}\frac{|\Pi|T}{\delta}},$
@@
$\widehat{V}(\pi_{\text{Dstl}};\mathcal{D}):=$ $\mathbb{E}_{\rho,\pi_{\text{Dstl}}}[{\widehat{r}}_{\text{Dstl}}]-\mathbb{E}_{\rho,\pi_{\texttt{ref}}}[{\widehat{r}}_{\text{Dstl}}]-\beta\text{KL}(\pi_{\text{Dstl}}\|\pi_{\texttt{ref}})$
$+\frac{1}{\eta}L_{\mathcal{D}}({\widehat{r}}_{\text{Dstl}})-\frac{1}{\eta}L_{\mathcal{D}}({\widehat{r}}_{\texttt{MLE}})-\texttt{bonus},$
@@
$\widehat{V}(\pi_{\text{Dstl}};\mathcal{D})$
$\leq$ $\mathbb{E}_{\rho,\pi_{\text{Dstl}}}[r^{*}]-\mathbb{E}_{\rho,\pi_{\texttt{ref}}}[r^{*}]-\beta\text{KL}(\pi_{\text{Dstl}}\|\pi_{\texttt{ref}})+\frac{1}{\eta}L_{\mathcal{D}}(r^{*})-\frac{1}{\eta}L_{\mathcal{D}}({\widehat{r}}_{\texttt{MLE}})-\texttt{bonus}$ (Pessimistic estimation of ${\widehat{r}}_{\text{Dstl}}$ in Eq. (12))
$\leq$ $J_{\beta}(\pi_{\text{Dstl}})-J_{\beta}(\pi_{\texttt{ref}})+\frac{2}{\eta|\mathcal{D}|}\log\frac{|\Pi|}{\delta}-\texttt{bonus}$ (Lem. H.2)
$\leq$ $J_{\beta}(\pi_{\text{Dstl}})-J_{\beta}(\pi_{\texttt{ref}})+2c\cdot e^{2{R}}\sqrt{\frac{1}{|\mathcal{D}|}\log\frac{|\Pi|T}{\delta}}-\texttt{bonus}.$ (14)
@@
$J_{\beta}(\pi)-J_{\beta}(\pi_{\texttt{ref}})-\widehat{V}(\pi_{\text{Dstl}};\mathcal{D})$
@@
$-\Big{(}\mathbb{E}_{\rho,\pi_{\text{Dstl}}}[{\widehat{r}}_{\text{Dstl}}]-\mathbb{E}_{\rho,\pi_{\texttt{ref}}}[{\widehat{r}}_{\text{Dstl}}]-\beta\text{KL}(\pi_{\text{Dstl}}\|\pi_{\texttt{ref}})+\frac{1}{\eta}L_{\mathcal{D}}({\widehat{r}}_{\text{Dstl}})\Big{)}+\frac{1}{\eta}L_{\mathcal{D}}({\widehat{r}}_{\texttt{MLE}})+\texttt{bonus}$
$\leq$ $\Big{(}\mathbb{E}_{\rho,\pi}[r^{*}]-\mathbb{E}_{\rho,\pi_{\texttt{ref}}}[r^{*}]-\beta\text{KL}(\pi\|\pi_{\texttt{ref}})\Big{)}-\min_{r\in\mathcal{R}^{\Pi}}\Big{(}\mathbb{E}_{\rho,\pi}[r]-\mathbb{E}_{\rho,\pi_{\texttt{ref}}}[r]-\beta\text{KL}(\pi\|\pi_{\texttt{ref}})+\frac{1}{\eta}L_{\mathcal{D}}(r)\Big{)}$ (Optimality of $\pi_{\text{Dstl}}$ in RPO;)
@@
$J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{\texttt{ref}})-\widehat{V}(\pi_{\text{Dstl}};\mathcal{D})\leq 4c_{2}\cdot e^{2{R}}\cdot{\texttt{Cov}}^{\pi^{*}_{r^{*}}|\pi_{\text{mix}}^{\mathcal{D}}}\cdot\sqrt{\frac{1}{|\mathcal{D}|}\log\frac{|\Pi|T}{\delta}}+\texttt{bonus}.$ (16)
@@
$J_{\beta}(\pi^{k-1}_{\text{mix}})-J_{\beta}(\pi_{\texttt{ref}})-\widehat{V}(\pi_{\text{Dstl}};\mathcal{D})\leq 4c_{2}\cdot e^{2{R}}\cdot\frac{2}{\alpha}\cdot\sqrt{\frac{1}{|\mathcal{D}|}\log\frac{|\Pi|T}{\delta}}.$
@@
$J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{\texttt{ref}})-\widehat{V}(\pi_{\text{Dstl}};\mathcal{D})$
$\leq$ $J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{{\text{mix}}}^{k-1})+J_{\beta}(\pi_{{\text{mix}}}^{k-1})-J_{\beta}(\pi_{\texttt{ref}})-\widehat{V}(\pi_{\text{Dstl}};\mathcal{D})$
@@
$\widehat{V}(\pi_{\text{Dstl}};\mathcal{D})\leq$ $J_{\beta}(\pi_{\text{Dstl}})-J_{\beta}(\pi_{\texttt{ref}})$
$\widehat{V}(\pi_{\text{Dstl}};\mathcal{D})\geq$ $J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{\texttt{ref}})-c^{\prime}{R}e^{2{R}}\cdot\Big{(}{\texttt{Cov}}^{\pi^{*}_{r^{*}}|\pi_{\text{mix}}^{\mathcal{D}}}\wedge\frac{\sqrt{\mathcal{C}(\Pi)}}{\alpha}\Big{)}\cdot\sqrt{\frac{1}{|\mathcal{D}|}\log^{c_{0}}\frac{|\Pi|T}{\delta}}.$